OpenStudy (anonymous):

A box of ground nutmeg weighs 1 1/3 ounce. If there are 20 teaspoons of nutmeg in the box, how much does one teaspoon of nutmeg weigh

OpenStudy (anonymous):

1/15 ounce 1/20 ounce 1/60 ounce 4/3 ounce

OpenStudy (anonymous):

these are the possible answers its giving me

OpenStudy (anonymous):

i would really appreciate it if i could get some help

OpenStudy (misty1212):

HI!!

OpenStudy (anonymous):

hi

OpenStudy (misty1212):

divide

OpenStudy (misty1212):

\[\huge 1\tfrac{1}{3}\div 20\]

OpenStudy (anonymous):

im getting 0.66

OpenStudy (misty1212):

turn the mixed number \(1\tfrac{1}{3}\) in to the improper fraction \(\frac{4}{3}\) and then do \[\huge \frac{4}{3}\times \frac{1}{20}\]

OpenStudy (misty1212):

your answer choices are not decimals, don't use a calculator

OpenStudy (misty1212):

cancel the 4 and get \[\huge \frac{1}{3}\times \frac{1}{5}\]

OpenStudy (anonymous):

so 1/15 would be the answer

OpenStudy (misty1212):

yes

OpenStudy (anonymous):

thank you so much :)

OpenStudy (misty1212):

\[\color\magenta \heartsuit\]